\begin{thebibliography}{10}

\bibitem{adams:1994}
W.~W. Adams and P.~Loustaunau.
\newblock {\em An introduction to {G}r\"obner bases}, volume~3 of {\em Graduate
  Studies in Mathematics}.
\newblock American Mathematical Society, Providence, RI, 1994.

\bibitem{bassa:2015}
A.~Bassa, P.~Beelen, A.~Garcia, and H.~Stichtenoth.
\newblock Towers of function fields over non-prime finite fields.
\newblock {\em Mosc. Math. J.}, 15(1):1--29, 181, 2015.

\bibitem{boonniyoma:2016}
K.~Boonniyoma and S.~Jitman.
\newblock Complementary dual subfield linear codes over finite fields.
\newblock {\em CoRR}, abs/1605.06827, 2016.

\bibitem{braun:2013}
M.~Braun, T.~Etzion, and A.~Vardy.
\newblock Linearity and complements in projective space.
\newblock {\em Linear Algebra Appl.}, 438(1):57--70, 2013.

\bibitem{bringer:2014}
J.~Bringer, C.~Carlet, H.~Chabanne, S.~Guilley, and H.~Maghrebi.
\newblock Orthogonal direct sum masking – a smartcard friendly computation
  paradigm in a code, with builtin protection against side-channel and fault
  attacks.
\newblock In {\em WISTP}, volume 8501 of {\em Lecture Notes in Comput. Sci.},
  pages 40--56. Springer, Berlin, 2014.

\bibitem{brylawski:1992}
T.~Brylawski and J.~Oxley.
\newblock The {T}utte polynomial and its applications.
\newblock In {\em Matroid applications}, volume~40 of {\em Encyclopedia Math.
  Appl.}, pages 123--225. Cambridge Univ. Press, Cambridge, 1992.

\bibitem{carlet:2014}
C.~Carlet and S.~Guilley.
\newblock Complementary dual codes for counter-measures to side-channel
  attacks.
\newblock In {\em Proceedings of the 4th ICMCTA Meeting}, volume~3 of {\em CIM
  Series in Mathematical Sciences book series}, pages 87--95. Springer, Berlin,
  2014.

\bibitem{carlet:2016}
C.~Carlet and S.~Guilley.
\newblock Complementary dual codes for counter-measures to side-channel
  attacks.
\newblock {\em Adv. Math. Commun.}, 10(1):131--150, 2016.

\bibitem{carlet:2017}
C.~Carlet, S.~Mesnager, C.~Tang, and Y.~Qi.
\newblock Euclidean and {H}ermitian {LCD} {MDS} codes.
\newblock {\em CoRR}, abs/1702.08033, 2017.

\bibitem{chen:2014}
B.~Chen, H.Q. Dinh, and H.~Liu.
\newblock Repeated-root constacyclic codes of length {$\ell p^s$} and their
  duals.
\newblock {\em Discrete Appl. Math.}, 177:60--70, 2014.

\bibitem{chen:2017}
B.~Chen and H.~Liu.
\newblock New constructions of {MDS} codes with complementary duals.
\newblock {\em CoRR}, abs/1702.07831, 2017.

\bibitem{cox:2005}
D.~A. Cox, J.~Little, and D.~O'Shea.
\newblock {\em Using algebraic geometry}, volume 185 of {\em Graduate Texts in
  Mathematics}.
\newblock Springer, New York, second edition, 2005.

\bibitem{cox:2015}
D.~A. Cox, J.~Little, and D.~O'Shea.
\newblock {\em Ideals, varieties, and algorithms}.
\newblock Undergraduate Texts in Mathematics. Springer, Cham, fourth edition,
  2015.
\newblock An introduction to computational algebraic geometry and commutative
  algebra.

\bibitem{dougherty:2015}
S.~T. {Dougherty}, J.-L. {Kim}, B.~{Ozkaya}, L.~{Sok}, and P.~{Sol{\'e}}.
\newblock The combinatorics of {LCD} codes: Linear programming bound and
  orthogonal matrices.
\newblock {\em International Journal of Information and Coding Theory},
  4(2-3):116–--128, 2015.

\bibitem{esmaeili:2009}
M.~Esmaeili and S.~Yari.
\newblock On complementary-dual quasi-cyclic codes.
\newblock {\em Finite Fields Appl.}, 15(3):375--386, 2009.

\bibitem{galvez:2017}
L.~Galvez, J.-L. Kim, N.~Lee, Y.~G. Roe, and B.-S. Won.
\newblock Some bounds on binary {LCD} codes.
\newblock {\em CoRR}, abs/1701.04165, 2017.

\bibitem{greene:1976}
C.~Greene.
\newblock Weight enumeration and the geometry of linear codes.
\newblock {\em Studies in Applied Mathematics}, 55:119--128, 1976.

\bibitem{gueneri:2016}
Cem G\"uneri, Buket \"Ozkaya, and Patrick Sol\'e.
\newblock Quasi-cyclic complementary dual codes.
\newblock {\em Finite Fields Appl.}, 42:67--80, 2016.

\bibitem{hoholdt:19998}
T.~H{\o}holdt.
\newblock On (or in) the {B}lahut footprint.
\newblock In {\em Codes, curves, and signals ({U}rbana, {IL}, 1997)}, volume
  485 of {\em Kluwer Internat. Ser. Engrg. Comput. Sci.}, pages 3--7. Kluwer
  Acad. Publ., Boston, MA, 1998.

\bibitem{hou:2016}
X.~Hou and F.~Oggier.
\newblock On {LCD} codes and lattices.
\newblock In {\em Proc. IEEE Int. Symp. on Inform. Theory 2016}, pages
  1501--1505. 2016.

\bibitem{jaeger:1989}
F.~Jaeger.
\newblock Tutte polynomials and bicycle dimension of ternary matroids.
\newblock {\em Proc. Amer. Math. Soc.}, 107(1):17--25, 1989.

\bibitem{jin:2017}
L~Jin.
\newblock Construction of {MDS} codes with complementary duals.
\newblock {\em IEEE Trans. Inform. Theory}, 63(5):2843--2847, 2017.

\bibitem{jurrius:2013}
R.~P. M.~J. Jurrius and R.~Pellikaan.
\newblock {\em Codes, arrangements and matroids}, pages 219--325.
\newblock Ser. Coding Theory Cryptol. World Scientific Publishing, 2013.

\bibitem{lang:1965}
Serge Lang.
\newblock {\em Algebra}.
\newblock Addison-Wesley Publishing Co., Inc., Reading, Mass., 1965.

\bibitem{leemans:2016}
D.~Leemans and B.~G. Rodrigues.
\newblock Linear codes with complementary duals from some strongly regular
  subgraphs of the {M}c{L}aughlin graph.
\newblock {\em Math. Commun.}, 21(2):239--249, 2016.

\bibitem{li:2017}
C.~Li.
\newblock On {H}ermitian {LCD} codes from cyclic codes and their applications
  to orthogonal direct sum masking.
\newblock {\em CoRR}, abs/1701.03986, 2017.

\bibitem{li:2016a}
C.~Li, C.~Ding, and S.~Li.
\newblock {LCD} cyclic codes over finite fields.
\newblock {\em CoRR}, abs/1608.02170, 2016.

\bibitem{li:2016b}
S.~Li, C.~Li, C.~Ding, and H.~Liu.
\newblock Parameters of two classes of {LCD} {BCH} codes.
\newblock {\em CoRR}, abs/1608.02670, 2016.

\bibitem{lina:2016}
E.~R. Lina~Jr. and E.~G. Nocon.
\newblock On the construction of some {LCD} codes over finite fields.
\newblock {\em Manila Journal of Science}, 9:67--82, 2016.

\bibitem{liu:2016}
X.~Liu and H.~Liu.
\newblock Matrix-product complementary dual codes.
\newblock {\em CoRR}, abs/1604.03774, 2016.

\bibitem{massey:1992}
J.L. Massey.
\newblock Linear codes with complementary duals.
\newblock {\em Discrete Math.}, 106/107:337--342, 1992.

\bibitem{mesnager:2016}
S.~Mesnager, C.~Tang, and Y.~Qi.
\newblock Complementary dual algebraic geometry codes.
\newblock {\em CoRR}, abs/1609.05649, 2016.

\bibitem{rosenstiehl:1978}
P.~Rosenstiehl and R.~C. Read.
\newblock On the principal edge tripartition of a graph.
\newblock {\em Ann. Discrete Math.}, 3:195--226, 1978.
\newblock Advances in graph theory (Cambridge Combinatorial Conf., Trinity
  College, Cambridge, 1977).

\bibitem{sangwisut:2015}
E.~Sangwisut, S.~Jitman, S.~Ling, and P.~Udomkavanich.
\newblock Hulls of cyclic and negacyclic codes over finite fields.
\newblock {\em Finite Fields Appl.}, 33:232--257, 2015.

\bibitem{sari:2016}
M.~Sari and M.~E. Koroglu.
\newblock On {MDS} negacyclic {LCD} codes.
\newblock {\em CoRR}, abs/1611.06371, 2016.

\bibitem{sendrier:1997}
N.~Sendrier.
\newblock On the dimension of the hull.
\newblock {\em SIAM J. Discrete Math.}, 10(2):282--293, 1997.

\bibitem{sendrier:2000}
N.~Sendrier.
\newblock Finding the permutation between equivalent linear codes: the support
  splitting algorithm.
\newblock {\em IEEE Trans. Inform. Theory}, 46(4):1193--1203, 2000.

\bibitem{sendrier:2002}
N.~Sendrier.
\newblock On the security of the {M}c{E}liece public-key cryptosystem.
\newblock In {\em Information, coding and mathematics}, volume 687 of {\em
  Kluwer Internat. Ser. Engrg. Comput. Sci.}, pages 141--163. Kluwer Acad.
  Publ., Boston, MA, 2002.

\bibitem{sendrier:2004}
N.~Sendrier.
\newblock Linear codes with complementary duals meet the {G}ilbert-{V}arshamov
  bound.
\newblock {\em Discrete Math.}, 285(1-3):345--347, 2004.

\bibitem{sendrier:2013a}
N.~Sendrier and D.E. Simos.
\newblock How easy is code equivalence over $\fq$?
\newblock In {\em WCC 2013, Apr 2013, Bergen, Norway}, International Workshop
  on Coding and Cryptography, pages 80--93. 2013.

\bibitem{sharma:2017}
A.~Sharma and T.~Kaur.
\newblock Enumeration of complementary-dual cyclic $\fq$-linear $\fqt$-codes.
\newblock {\em arXiv}, abs/1702.00549, 2017.

\bibitem{simonis:1994}
Juriaan Simonis.
\newblock The effective length of subcodes.
\newblock {\em Appl. Algebra Engrg. Comm. Comput.}, 5(6):371--377, 1994.

\bibitem{sok:2017}
L.~Sok, M.~Shi, and P.~Sol{\'{e}}.
\newblock Constructions of optimal {LCD} codes over large finite fields.
\newblock {\em CoRR}, abs/1704.04083, 2017.

\bibitem{tsfasman:1982}
M.~A. Tsfasman, S.~G. Vl\v{a}du\c{t}, and Th. Zink.
\newblock Modular curves, {S}himura curves, and {G}oppa codes, better than
  {V}arshamov-{G}ilbert bound.
\newblock {\em Math. Nachr.}, 109:21--28, 1982.

\bibitem{kandasamy:2012}
W.~B. Vasantha~Kandasamy, F.~Smarandache, R.~Sujatha, and R.~S. Raja~Duray.
\newblock {\em Erasure Techniques in {MRD} Codes}.
\newblock ZIP Publishing, Ohio, 2012.

\bibitem{yang:1994}
X.~Yang and J.L. Massey.
\newblock The condition for a cyclic code to have a complementary dual.
\newblock {\em Discrete Math.}, 126(1-3):391--393, 1994.

\end{thebibliography}

\documentclass[a4wide,12pt]{article}

\usepackage[pdftex]{graphicx}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage[textsize=small,textwidth=1.3in]{todonotes}
\usepackage[english]{babel}

\usepackage{hyperref}

\newtheorem{prop}{Proposition}[section]
\newtheorem{thm}[prop]{Theorem}

\newtheorem{cor}[prop]{Corollary}
\newtheorem{Defi}[prop]{Definition}
\newtheorem{Rem}[prop]{Remark}
\newtheorem{Exam}[prop]{Example}
\newtheorem{Exer}[prop]{Exercise}
\newtheorem{Expe}[prop]{Experiment}
\newtheorem{Cons}[prop]{Construction}

\newtheorem{Alg}[prop]{Algorithm}
\newtheorem{Prob}[prop]{Research Problem}
\newenvironment{defi}{\begin{Defi} \rm}{\end{Defi}}
\newenvironment{rem}{\begin{Rem} \rm}{\end{Rem}}
\newenvironment{ex}{\begin{Exam} \rm}{\end{Exam}}

\newenvironment{proof}{\noindent{\bf Proof.}\ }{\hspace*{\fill}$\diamond$\medskip\par}

\newcommand{\fq}{{\mathbb F}_q}

\newcommand{\fqt}{{\mathbb F}_{q^t}}

\newcommand{\ba}{{\bf a}}
\newcommand{\bb}{{\bf b}}
\newcommand{\bc}{{\bf c}}

\newcommand{\bg}{{\bf g}}

\newcommand{\bx}{{\bf x}}
\newcommand{\by}{{\bf y}}

\title{LCD codes over ${\mathbb F}_q $ are as good\\
 as linear codes for $q$ at least four}

\author{Ruud Pellikaan
\footnote{Department of Mathematics and Computing Science, Eindhoven University of Technology, P.O. Box 513, 5600 MB  Eindhoven, The Netherlands. E-mail: g.r.pellikaan@tue.nl}}

\begin{document}

\maketitle

\begin{abstract}
\noindent The hull $H(C)$ of a linear code $C$ is defined by $H(C)=C \cap C^\perp$.
A linear code with a complementary dual (LCD) is a linear code with $H(C)=\{0\}$.
The dimension of the hull of a code is an invariant under permutation equivalence.
For binary and ternary codes the dimension of the hull is also invariant under monomial equivalence
and we show that this invariant is determined by the extended weight enumerator of the code.\\
The hull of a code is not invariant under monomial equivalence if $q\geq 4$.
We show that every ${\mathbb F}_q $-linear code is monomial equivalent with an LCD code in case $q \geq 4$.
The proof uses techniques from Gr\"obner basis theory.
We conclude that if there exists an ${\mathbb F}_q $-linear code with parameters $[n,k,d]_q$ and $q \geq 4$,
then there exists also a LCD code with the same parameters.
Hence this holds for optimal and MDS codes.
In particular there exist LCD codes that are above the Gilbert-Varshamov bound if $q$ is a square and $q\geq 49$
by the existence of such codes that are algebraic geometric.\\
Similar results are obtained with respect to Hermitian LCD codes.
\end{abstract}

\newpage

\section{Introduction}

\noindent The hull $H(C)$ of a linear code $C$ is defined by $H(C)=C \cap C^\perp$.
A linear code with a complementary dual (LCD) is a linear code with $H(C)=\{0\}$.
LCD codes were introduced and studied by Massey \cite{massey:1992} who showed that
these codes are optimal for the two-user binary adder channel (BAC) and that they are asymptotically good.
Sendrier \cite{sendrier:2004} showed that they meet the Gilbert-Varshamov bound and he
\cite{sendrier:1997,sendrier:2000,sendrier:2004,sendrier:2013a} considered the hull of a code
to find a permutation between two equivalent codes
with an application to code-based public key cryptosystems  \cite{sendrier:2002}. \\
Carlet and Guilley gave applications of LCD codes in side-channel attacks (SCA) and fault non-invasive attacks \cite{bringer:2014,carlet:2014,carlet:2016,li:2017},
and may be used for the construction of lattices \cite{hou:2016}. LCD rank metric codes have applications in network coding \cite{braun:2013,kandasamy:2012}.\\
Constructions  of LCD codes are known for the following types of codes:
cyclic \cite{li:2016a,li:2016b,sangwisut:2015,yang:1994}, consta-cyclic \cite{chen:2014,sari:2016} and quasi-cyclic \cite{esmaeili:2009,gueneri:2016},
graphic \cite{leemans:2016} and algebraic geometric \cite{mesnager:2016}.
Optimal and MDS codes that are LCD are considered in
\cite{boonniyoma:2016,carlet:2017,chen:2017,dougherty:2015,galvez:2017,jin:2017,lina:2016,liu:2016,sari:2016,sok:2017}.\\
Apart from the standard or Euclidean inner product,
the notion of LCD codes is researched with respect to the Hermitian inner product in case $q$ is a square \cite{boonniyoma:2016,carlet:2017,li:2017,sharma:2017}.\\
The dimension of the hull of a code is an invariant under permutation equivalence.
For binary and ternary codes the dimension of the hull is also invariant under monomial equivalence
and we show that this invariant is determined by the extended weight enumerator of the code.\\
The hull of a code is not invariant under monomial equivalence if $q\geq 4$.\\
We show that every ${\mathbb F}_q $-linear code is monomial equivalent with an LCD code in case $q \geq 4$,
by means of the theory of Gr\"{o}ber bases generalizing a technique in \cite{carlet:2017}.
We conclude that if there exists an ${\mathbb F}_q $-linear code with parameters $[n,k,d]_q$ and $q \geq 4$,
then there exists also a LCD code with the same parameters.
In particular this holds for optimal and MDS codes.
And there exist LCD codes that are above the Gilbert-Varshamov bound if $q$ is a square and $q\geq 49$
by the existence of such codes that are algebraic geometric.\\
Similar results are obtained with respect to Hermitian LCD codes.

\section{Preliminaries}

\noindent
Let  $q $ be a power of a prime. Then $\fq$ denotes the finite field with $q$ elements.\\
An $[n,k,d]_q$ code is a $k$-dimensional subspace of $\fq^n$ of minimum distance $d$.

\begin{defi}
The {\em standard} or {\em Euclidean inner product} of $\ba$ and $\bb$ in $\fq^n$ is defined by
$$
\ba\cdot \bb = \sum_{i=1}^n a_ib_i.
$$
Now $A \perp B$ if and only if $\ba \cdot \bb = 0$ for all $\ba \in A$ and $\bb \in B$.
The {\em dual} of $C$ is defined by
$$
C^\perp = \{ \ \bx \in \fq^n \ | \ \bx\cdot \bc =0 \mbox{ for all } \bc \in C \ \}.
$$
Let $q$ be an even power of a prime. Define the conjugate $\bar{x}$ of $x \in \fq$ by  $\bar{x}=x^{\sqrt{q}}$.
The {\em Hermitian inner product} of $\ba$ and $\bb$ in $\fq^n$ is defined by
$$
\ba \cdot_H \bb = \sum_{i=1}^n a_i\bar{b}_i.
$$
And $A \perp_H B$ if and only if $\ba \cdot_H \bb = 0$ for all $\ba \in A$ and $\bb \in B$.
The {\em Hermitian dual} of $C$ is defined by
$$
C^{\perp_H} = \{ \ \bx \in \fq^n \ | \ \bx\cdot_H \bc =0 \mbox{ for all } \bc \in C \ \}.
$$
\end{defi}

\begin{defi}Let $C$ be an $\fq $-linear code. \\
The {\em hull} of $C$ is defined by $H(C)=C \cap C^\perp $.
If $H(C) = \{ {\bf 0 } \} $, then $C$ is called a {\em linear} code with {\em complementary dual} or an LCD code.\\
The code $C$ is called {\em Hermitian} LCD if $C \cap C^{\perp_H} = \{ 0 \}$.
\end{defi}

\begin{defi}A {\em permutation} matrix  is a square matrix with zeros and ones, such that in every
row (and in every column) there is exactly one element equal to one.
A {\em diagonal} matrix is a square matrix with zeros outside its diagonal.
A {\em monomial} matrix  is a square matrix such that in every
row (and in every column) there is exactly one nonzero element.
\end{defi}

\begin{rem}A permutation matrix and an invertible diagonal matrix are special monomial matrices.
\end{rem}

\begin{defi}Let $C_1$ and $C_2$ be $\fq $-linear codes in $\fq^n$.
Then $C_1$ is called {\em equivalent} or {\em permutational equivalent} to $C_2$,
and is denoted by $C_1\equiv C_2$ if there exists a permutation matrix $\Pi$ such that $\Pi(C_1)=C_2$.
The code $C_1$ is {\em scalar equivalent} with $C_2$ if there is an invertible diagonal matrix $D$
such that $D(C_1)=C_2$.
And $C_1$ is called {\em monomial equivalent} to $C_2$,
denoted by $C_1\cong C_2$ if there exists a monomial matrix $M$ such that $M(C_1)=C_2$.
\end{defi}

\begin{defi}Let $\bx$, $\by  \in \fq ^n$.
Then the {\em star product} is defined by
$$
\bx *\by =(x_1y_1,\ldots ,x_ny_n).
$$
Let $\bx \in \fq ^n$ with nonzero entries, then $\bx^{-1}=(x_1^{-1},\ldots ,x_n^{-1})$.\\
Let $C \subseteq \fq^n$. Then $\bx *C = \{ \ \bx *\bc \ | \ \bc \in C \ \}$.\\
Let $G$ be an $k\times n$ matrix with $i$-th row $\bg_i$.
Then $\bx*G$ is the $k\times n$ matrix with $i$-th row $\bx*\bg_i$.
\end{defi}

\begin{rem}Note that $C_1$ and $C_2$ are scalar equivalent if and only if
there exists an $\bx $ with nonzero entries such that $C_2=\bx * C_1$.\\
Let $C$ be an $\fq$-linear code of length $n$ and $\bx \in \fq ^n$ with nonzero entries.\\
If $G$ is a generator matrix and $H$ a parity check matrix of $C$,
then $\bx*G$ is a generator matrix and $\bx ^{-1}*H$ is a parity check matrix of $\bx*C$.
\end{rem}

\section{The dimension of the hull of a code}
The dimension of the hull of a code is considered in \cite{sendrier:1997}.\\
The dimension of $H(C)$ is an invariant of permutational equivalent codes but not for monomial equivalent codes,
except for $q=2$ and $q=3$, where it can be expressed in terms of the Tutte polynomial of the matroid of $C$.\\
Let $M(C)$ be the matroid of the code $C$ and $t_{M(C)}(X,Y)$ the two variable Tutte polynomial of the matroid.
For a binary code $C$ we have
$$
|t_{M(C)}(-1,-1)|= 2^{\dim H(C)}.
$$
See \cite{rosenstiehl:1978} for graph codes and \cite[Proposition 6.5.4]{brylawski:1992} for arbitrary codes.\\
For a ternary code $C$ we have
$$
|t_{M(C)}(j,j^2)| = (\sqrt{3})^{\dim H(C)},
$$
where $j=e^{2\pi i/3}$ is a primitive third root of unity of the complex numbers and $|z|$ is the modulus of the complex number $z$. See \cite{jaeger:1989}.\\
$W_C(X,Y)$, the weight enumerator  of an $\fq $-linear code $C$ can be expressed in terms of $t_{M(C)}(X,Y)$ as follows:
\[ W_C(X,Y)=(X-Y)^kY^{n-k}\ t_C\left(\frac{X+(q-1)Y}{X-Y},\frac{X}{Y}\right)\; . \]
See Greene \cite{greene:1976}. But the converse is not true.

\begin{ex}\label{ex-1}
The dimension of the hull of a binary code $C$ cannot be expressed in terms of $W_C(X,Y)$.
Consider the two codes $C_1$ and $C_2$ of \cite[Example 1]{simonis:1994} with generator matrices
$G_1=(I_3|I_3)$ and $G_2=(I_3|J_3)$, where $J_3$ is the $3\times 3$ all ones matrix.
Both have the same weight enumerator:
$$
W_{C_1}(X,Y)=W_{C_2}(X,Y)=X^6+3X^4Y^2+3X^2Y^4+Y^6.
$$
Furthermore $C_1$ is self dual, so $\dim H(C_1)=3$.
But $H(C_2)$ is generated by the all ones vector, so $\dim H(C_2)=1$.
\end{ex}

\begin{rem}
Let $t_{C}(X,Y)=t_{M(C)}(X,Y)$. Then
$t_{C}(X,Y)$ can be expressed in $W_C(X,Y,T)$, the extended weight enumerator  of $C$ and vice versa. See \cite[Theorems 5.9 and 5.10]{jurrius:2013}.
\[ W_C(X,Y,T)=(X-Y)^kY^{n-k}\ t_C\left(\frac{X+(T-1)Y}{X-Y},\frac{X}{Y}\right)\; . \]
\[ t_C(X,Y)=Y^n(Y-1)^{-k}W_C(1,Y^{-1},(X-1)(Y-1))\; . \]
So we have in the binary case
\[ |W_C(1,-1,4)|=2^k|t_C(-1,-1)|=2^{k+h}. \]
where $k$ and $h$ are the dimensions of $C$ and $H(C)$, respectively.
\end{rem}

\begin{ex}\label{ex-2}For the extended weight enumerator $W_C(X,Y,T)-X^n $ is divisible by $T-1$.
Define $\bar{W}_C(X,Y,T)=(W_C(X,Y,T)-X^n)/(T-1)$.\\
In the previous Example \ref{ex-1} we have that
$$
\bar{W}_{C_1}(X,Y,T)=3X^4Y^2+3(T-1)X^2Y^4+(T-1)^2Y^6 
$$
So $W_{C_1}(1,-1,4)=2^6$. And $\bar{W}_{C_2}(X,Y,T)$ is equal to
$$
3X^4Y^2+(T-2)X^3Y^3+3X^2Y^4+3(T-2)XY^5+(T^2-3T+3)Y^6.
$$
So $W_{C_2}(1,-1,4)=2^4$.
\end{ex}

\begin{rem}
For the ternary case  we have
\[ t_C(j,j^2)=
j^{2n}(j^2-1)^{-k}W_C(1,j) \; , \]
since $j^{-2}=j$ and $(j-1)(j^2-1)=3$.
So the dimension of the hull of a ternary code $C$ can be expressed in terms of $W_C(X,Y)$.
Furthermore
\[ W_C(1,j)= a_0 + a_1j + a_2j^2\; , \]
where $a_0=\sum_{i=0}^{n/3} A_{3i}$, $a_1=\sum_{i=0}^{n/3} A_{3i+1}$ and $a_2=\sum_{i=0}^{n/3} A_{3i+2}$.
\end{rem}

\section{Existence of LCD codes}

\begin{prop}\label{prop-GGT}Let $C$ be an $\fq$-linear $[n,k]$ code with generator matrix $G$.
Let $h$ be the dimension of $H(C)$ and $r=k-h$. Then $C$ has a generator matrix $G_0$ such that
$$
G_0G_0^T=
\left(
\begin{array}{c|c}
O_{h\times h}&O_{h\times r} \\
\hline
O_{r \times h} &P \\
\end{array}
\right),
$$
where $O_{l\times m}$ is the all zeros $l\times m$ matrix and $P$ is an invertible $r \times r$ matrix.
Furthermore the rank of $GG^T$ is $r$ for every generator matrix $G$ of $C$.
\end{prop}

\begin{proof}Let $\bg_1, \ldots \bg_h$ be a basis of $H(C)$. Now $H(C)$ is a sub space of $C$,
so we can extend $\bg_1, \ldots \bg_h$ to a basis $\bg_1, \ldots \bg_k$ of $C$.
Let $G_0$ be the $k \times n$ matrix with $i$-th row $\bg_i$. Then $G_0$ is a generator matrix of $C$ and
$G_0G_0^T$ is a $k \times k$ matrix with $\bg_i \cdot \bg_j$ at position $(i,j)$.
Now  $\bg_i \cdot \bg_j=0$ if $i\leq h$ or $j\leq h$,
since $\bg_i, \bg_j \in C$ for all $1 \leq i,j \leq n$ and $\bg_i \in C^\perp$ if $i\leq h$ and $\bg_j \in C^\perp$ if $j\leq h$.
Therefore $G_0G_0^T$ has the form as stated in the Proposition.\\
Now suppose that $P$ is not invertible. Then the rows of $P$ are dependent and
there exists an invertible $r\times r$ matrix $M$ such that
the first row of $MP$ consists of zeros. But then also the the first row of $MPM^T$ consists of zeros.
Let $G_f$ be the submatrix of $G$ consisting of the first $h$ rows of $G$,
and $G_l$ the submatrix of $G$ consisting of the last $r$ rows of $G$.
Let $G_1$ be the $k \times n$ matrix with $G_f$ in the first $h$ rows and $MG_l$ in the last $r$ rows.
Then $G_1$ is also a generator matrix of $C$, since $M$ is invertible. Moreover
$$
G_1G_1^T=
\left(
\begin{array}{c|c}
O_{h\times h}&O_{h\times r} \\
\hline
O_{r \times h} &MPM^T \\
\end{array}
\right).
$$
Let $\bg'_i$ be the $i$-th row of $G_1$. Then $\bg'_i \cdot \bg'_j=0$ if $i\leq h+1$ and $j\leq k$,
since the first row of $MPM^T$ consists of zeroes. Hence $H(C)$ consists of $h+1$ independent elements
$\bg'_1, \ldots \bg'_h, \bg'_{h+1}$ which is a contradiction.
Therefore $P$ is invertible and $G_0G_0^T$ has rank $r$.\\
Now suppose that $G$ is a generator matrix of $C$. Then there exists an invertible $k \times k$ matrix $N$
such that $G=NG_0$. Hence $GG^T=NG_0G_0^TN^T$. The rank of $G_0G_0^T$ is $r$ and this rank does not change
by elementary row operations, that is by multiplication on the left by $N$
and also not by elementary column operations, that is by multiplication on the right by $N^T$.
Therefore $GG^T$ has also rank $r$.\\
Remark that with the theory of symmetric forms one can show a more restrictive normal form of $P$. See \cite[Chap. XIV]{lang:1965}.
\end{proof}

\begin{cor}\label{cor-GGT}Let $C$ be an $\fq$-linear $[n,k]$ code with generator matrix $G$.\\
Then $C$ is an LCD code if and only if $GG^T$ is nonsingular. 
\end{cor}

\begin{proof}This is \cite[Proposition 1]{massey:1992} and a consequence of Proposition \ref{prop-GGT}.
\end{proof}

\begin{cor}\label{cor-HGGT}Let $C$ be an $\fq$-linear $[n,k]$ code with generator matrix $G$.\\
Then $C$ is a Hermitian LCD code if and only if $G\bar{G}^T$ is nonsingular. 
\end{cor}

\begin{proof}This is \cite[Theorem 3.4]{boonniyoma:2016}.
\end{proof}

\noindent
The multivariate notation is used for polynomials, that means that $X$ is an abbreviation of $(X_1, \ldots , X_n)$ and
the polynomial $f(X)$ of $f(X_1, \ldots , X_n)$.

\begin{prop}Let $f(X)$ be a nonzero polynomial of $\fq [X_1,\ldots ,X_n]$
such that the degree of $f(X)$ with respect to $X_j$ is at most $q-1$ for all $j$.
Then there exists a $\bx \in \fq ^n$ such that $f(\bx)\not=0$.
\end{prop}

\begin{proof}See \cite[V, \S 4, Theorem 5]{lang:1965}.
\end{proof}

\begin{prop}\label{prop-fX=0}Let $f(X)$ be a nonzero polynomial of $\fq [X_1,\ldots ,X_n]$
such that the degree of $f(X)$ with respect to $X_j$ is at most $q-2$ for all $j$.
Then there exists a $ \bx \in (\fq \setminus \{ 0 \}) ^n$ such that $f (\bx )\not=0$.
\end{prop}

\begin{proof}For the theory of Gr\"{o}bner bases and the concepts of a footprint and delta set we refer to  \cite{adams:1994,cox:2005,cox:2015,hoholdt:19998}.\\\
Let $f(X)$ be a nonzero polynomial of $\fq [X_1,\ldots ,X_n]$
such that the degree of $f(X)$ with respect to $X_j$ is at most $q-2$ for all $j$.\\
Consider the ideal $I_{q,n}$ in $ \fq [X_1,\ldots ,X_n]$ generated by the $X_j^{q-1}-1$, $j=1, \ldots ,n$.
Then $(\fq \setminus \{ 0 \}) ^n$ is the zero set of $I_{q,n}$, and $I_{q,n}$
is the vanishing ideal of $(\fq \setminus \{ 0 \}) ^n$.
The footprint or delta set of the $X_j^{q-1}-1$, $j=1, \ldots ,n$ is equal to $(\{ 0, 1, \ldots ,q-2\})^n$ and contains
the delta set of the ideal of  $I_{q,n}$. The delta set of $I_{q,n}$ is finite and has size $(q-1)^n$,
the size of the zero set of $I_{q,n}$ and it is equal to the size of delta set of the $X_j^{q-1}-1$, $j=1, \ldots ,n$ .
Hence the delta set of $I_{q,n}$ is equal to $(\{ 0, 1, \ldots ,q-2\})^n$ and  $\{ X_j^{q-1}-1 | j=1, \ldots ,n \}$ is a
Gr\"obner basis of $I_{q,n}$ with respect tot the total degree lex order. \\
Now $f(X)$ is a nonzero element of $\fq [X_1,\ldots ,X_n]$.
The exponent of the leading monomial of $f(X)$ is in the delta set of $I_{q,n}$,
since the degree of $f(X)$ with respect to $X_j$ is at most $q-2$ for all $j$.
Therefore $ f(X)\not\in I_{q,n}$.\\
Suppose that $f (\bx ) =0$ for all $ \bx \in (\fq \setminus \{ 0 \}) ^n$.
Then $f(X)$ is in the vanishing ideal of $(\fq \setminus \{ 0 \}) ^n$, which is $I_{q,n}$. So $ f(X)\in I_{q,n}$. This is a contradiction.\\
Therefore there exists a $ \bx \in (\fq \setminus \{ 0 \}) ^n$ such that $f (\bx )\not=0$.
\end{proof}

\begin{thm}\label{thm-LCD}Let $C$ be an $\fq$-linear code with $q\geq 4$. Then $C$ is monomial equivalent with an LCD code.
\end{thm}

\begin{proof}Let $C$ be an $\fq$-linear code with $q\geq 4$. \\
Without loss of generality we may assume that $C$ has a generator matrix of the form $G=(I_k|B)$.
Let $\bx =(x_1, \ldots , x_k)$ be an $k$-tuple of nonzero elements of $\fq $.
Let $D(\bx)$ be the diagonal matrix with $\bx$ on its diagonal.
Let $(D(\bx)|B)$ be the generator matrix of the code $C_{\bx}$. Then $C_{\bx}$ is monomial equivalent with $C$.
Let $X=(X_1, \ldots , X_k)$. Now define
$$
f(X)=\det ((D(X)|B)(D(X)|B)^T).
$$
Then
$$
f(X) = \det (D(X_1^2, \ldots , X_k^2)+BB^T)
$$
Hence $f(X)$ is a polynomial in the variables $X_1, \ldots , X_k$ and the degree
of $f(X)$ with respect to $X _i$ is $2$ for all $i$, which is at most $q-2$, since $q\geq 4$.
The leading term of $f(X)$ with respect to the total degree lex order is $X_1^2 \cdots X_k^2$.
So $f(X)$ is a nonzero polynomial. Therefore $f(\bx)\not=0$ for some
$\bx \in (\fq \setminus \{ 0 \}) ^k$ by Proposition \ref{prop-fX=0}.  \\
Hence $C_{\bx}$ is an LCD code for this choice of $\bx$ by Corollary \ref{cor-GGT}.
\end{proof}

\begin{thm}\label{thm-HLCD}Let $C$ be an $\fq$-linear code with $q$ a square and $q> 4$.
Then $C$ is monomial equivalent with a Hermitian LCD code.
\end{thm}

\begin{proof}The proof is similar to the proof of Theorem \ref{thm-LCD}.
Let $C$ be an $\fq$-linear code with $q$ a square and $q> 4$. \\
Without loss of generality we may assume that $C$ has a generator matrix of the form $G=(I_k|B)$.
Let $\bx =(x_1, \ldots , x_k)$ be an $k$-tuple of nonzero elements of $\fq $.
Let $D(\bx)$ be the diagonal matrix with $\bx$ on its diagonal.
Let $(D(\bx)|B)$ be the generator matrix of the code $C_{\bx}$. Then $C_{\bx}$ is monomial equivalent with $C$.
Let $X=(X_1, \ldots , X_k)$. Now define
$$
g(X)=\det ((D(X)|B)(D(X^{\sqrt{q}})|\bar{B})^T).
$$
Then
$$
g(X) = \det (D(X_1^{\sqrt{q}+1}, \ldots , X_k^{\sqrt{q}+1})+B\bar{B}^T)
$$
Hence $g(X)$ is a polynomial in the variables $X_1, \ldots , X_k$ and the degree
of $g(X)$ with respect to $X _i$ is $\sqrt{q}+1$ for all $i$, which is at most $q-2$, since $q> 4$.
The leading term of $f(X)$ with respect to the total degree lex order is $X_1^{\sqrt{q}+1} \cdots X_k^{\sqrt{q}+1}$.
So $f(X)$ is a nonzero polynomial. Therefore $f(\bx)\not=0$ for some
$\bx \in (\fq \setminus \{ 0 \}) ^k$ by Proposition \ref{prop-fX=0}.  \\
Hence $C_{\bx}$ is a Hermitian LCD code for this choice of $\bx$ by Corollary \ref{cor-HGGT}.
\end{proof}

\begin{rem}Both Theorems \ref{thm-LCD} and \ref{thm-HLCD} are generalizations of \cite[Proposition 3.1]{carlet:2017} and \cite[Proposition 4.1]{carlet:2017}, respectively.
As a result one can now state that the existence of an MDS code with given parameters $[n,k,n-k+1]_q$ is equivalent to the existence of a LCD MDS code with the same parameters if $q\geq 4$,
and to the existence of  Hermitian LCD MDS codes if $q > 4$.
And similar statements hold for optimal codes. See
\cite{boonniyoma:2016,carlet:2017,chen:2017,dougherty:2015,galvez:2017,jin:2017,lina:2016,liu:2016,sari:2016,sok:2017}.
\end{rem}

\begin{rem}There exist (Hermitian) LCD codes that are above the Gilbert-Varshamov (GV) bound if $q$ is a square and $q\geq 49$, by Theorem \ref{thm-LCD}
and since there are algebraic geometric codes above the GV bound. See \cite{bassa:2015,tsfasman:1982}.
\end{rem}

\section{Conclusion}
The dimension of the hull of a code is determined by the extended weight enumerator of the code in case $q=2$ or $q=3$.
We show that every ${\mathbb F}_q $-linear code is monomial equivalent with an LCD code in case $q \geq 4$,
and similarly with an Hermitian LCD in case $q$ is an even power of a prime and $q > 4$.

\bibliographystyle{plain}

\end{document}